# From paintbrush to pixel: A review of deep neural networks in AI-generated art

### v2

@@ 2.4 Transformers
Convolution has become the central component for image processing applications as neural networks progressed throughout time. However, the human visual system has evolved to be sparse and efficient. Therefore, we do not process our entire visual field with the same resources, unlike convolutions. Rather, our eyes perform a fixation point strategy by means of a visual attention system which plays an essential role in human cognition Noton and Stark (1971a, b). Inspired by the attention system in human vision, computational attention mechanisms have been developed and integrated into neural networks. One of the goals is reducing the computational burden caused by the convolution operation, while still improving the performance.
@@
These attention-based neural networks have been applied to a variety of applications such as image recognition or object tracking; see Soydaner (2022) for a review. In particular, a novel attention-based encoder-decoder neural network architecture was presented for neural machine translation (NMT) in 2015 Bahdanau et al. (2015). The idea behind this approach is illustrated in Fig. 5, which shows how the attention mechanisms in neural networks work. In this example, the encoder takes an input sentence in English, and the decoder outputs its translation in Dutch. Both encoder and decoder include recurrent neural networks (RNNs); see Graves (2012) for more information about RNNs. Basically, the encoder outputs hidden states, and the decoder takes all the hidden states as inputs. Before processing them, the decoder applies an attention mechanism that gives each hidden state a score. Then, it multiplies each hidden state by its score to which a softmax function is applied. The weighted scores are summed, and the result leads to the context vector for the decoder. By obtaining weighted hidden states which are most associated with certain words, the decoder focuses on the relevant parts of the input during decoding.
@@
Fig. 5: A neural machine translation example. The model takes an English sentence as input and translates it into Dutch. The figure shows encoder hidden states, and which words the model focuses more on (indicated by the color intensity) while translating.
@@
Attention mechanisms in neural networks have progressed rapidly, especially for NMT. One is the self-attention mechanism, which is the core building block of the Transformer Vaswani et al. (2017). The original Transformer is composed of encoder-decoder stacks which are entirely based on self-attention without any convolutional or recurrent layers. There are six identical layers in each of the encoder-decoder stacks that form the Transformer. To illustrate the model, only one encoder-decoder stack is shown in Fig. 6.
@@
Fig. 6: The Transformer architecture in detail Vaswani et al. (2017); Soydaner (2022). (Left) The Transformer with one encoder-decoder stack. (Center) Multi-head attention. (Right) Scaled dot-product attention.
@@
The encoder-decoder stacks in the Transformer are composed of fully-connected layers and multi-head attention which is a kind of self-attention mechanism applying scaled dot-product attention within itself. As seen in Fig. 6, these attention mechanisms use three vectors for each word, namely query (Q), key (K) and value (V). These vectors are computed by multiplying the input with weight matrices $W_{q}$, $W_{k}$ and $W_{v}$ which are learned during training. In general, each value is weighted by a function of the query with the corresponding key. The output is computed as a weighted sum of the values. In the scaled dot-product attention, the dot products of the query with all keys are computed. As given in Eq. 6, each result is divided by the square root of the dimension of the keys (to have more stable gradients) and passed onto the softmax function. Finally, each softmax score is multiplied with the value Vaswani et al. (2017):
@@
Multi-head attention extends this idea by applying linear activations to the inputs (keys, values and queries) h times based on different, learned linear representations (Fig. 6). Each of the projected versions of queries, keys and values are called heads, for which the scaled dot-product attention is performed in parallel. Thus, the self-attention is calculated multiple times using different sets of query, key and value vectors. This allows the model to jointly attend to information at different positions Vaswani et al. (2017). In the last step, the projections are concatenated. Additionally, the decoder applies masked multi-head attention to ensure that only previous word embeddings (tokens) are used when predicting the next word in the sentence.
There are different Transformer architectures for various tasks Khan et al. (2022); Lin et al. (2022). After the Transformer yielded major progress in natural language processing (NLP), it has been adapted to image processing tasks. Image Transformer Parmar et al. (2018) is one of these adaptations in which the Transformer is applied to image processing. Image Transformer applies self-attention in local neighborhoods for each query pixel, and performs well on image generation and image super-resolution. However, the current state-of-the-art model is Vision Transformer (ViT) Dosovitskiy et al. (2021), which splits an input image into patches before the Transformer takes the linear embeddings of these patches in sequence as input (Fig. 7). Vision Transformer performs well on image classification tasks whilst using fewer computational resources. Recent models, such as Swin transformer Liu et al. (2021), DeiT Touvron et al. (2021), BEiT Bao et al. (2021), and DINOv2 Oquab et al. (2024), have built on ViT and have been successful in image processing tasks.
Fig. 7: The Vision Transformer Dosovitskiy et al. (2021). In order to classify an image, it takes the input as patches, projects linearly, adds position embeddings, and uses a Transformer encoder.
@@ 2.5 Diffusion models
Today, text-to-image models such as DALL-E 3 Betker et al. (2023) or Midjourney have turned AI into a popular tool to produce mesmerizing images. These are diffusion models which have shown great success in generating high-quality images. They have shown to outperform GANs at image synthesis Dhariwal and Nichol (2021).
@@
Unlike GANs, the training of diffusion models does not require an adversarial setting. The original Denoising Diffusion method Sohl-Dickstein et al. (2015) was inspired by non-equilibrium thermodynamics that systematically destroys structure in a data distribution, then restores the data. Based on this method, denoising diffusion probabilistic models or diffusion models in short, have been applied to image generation Ho et al. (2020).
@@
Diffusion models require two main steps in the training phase (Fig. 8). At first, during the forward (diffusion) process, random noise is gradually added to the input image until the original input becomes all noise. This is performed by a fixed Markov chain which adds Gaussian noise for T successive steps. Secondly, during the reconstruction or reverse process, the model reconstructs the original data from the noise obtained in the forward process. The reverse process is defined as a Markov Chain with learned Gaussian transitions. Accordingly, the prediction of probability density at time t depends only on the state attained at time (t-1). Here, $x_{1},...,x_{T}$ are latents of the same dimensionality as the data which make the diffusion models latent variable models Ho et al. (2020).
@@
Fig. 8: The working principle of diffusion model in general Ho et al. (2020). Starting from an image, the forward process involves gradually adding noise (following a fixed Markov chain) until the image is all noise. In the reverse process, the original image is reconstructed step by step through a learned Markov Chain.
@@
The general structure of a diffusion model is given in Fig. 8. Estimating the probability density at an earlier time step, given the current state of the system, is non trivial, so the reverse process requires training a neural network. To this end, all previous gradients are necessary for obtaining the required estimation. For each step in the Markov Chain, a neural network learns to denoise the image. Optionally, the denoising process can be guided by text (see Section 3.3.2). In this case, a Transformer encoder maps a text prompt to tokens which are then fed to the neural network (Fig. 9). Once trained, a diffusion model can be used to generate data by simply passing random noise (and optionally a text prompt) through the learned denoising process. It should be noted that this process requires multiple denoising steps at inference time, making diffusion models slow in generating images when prompted. As a result, several methods have been proposed to speed up inference from multi-step to few-step and even single-step Song et al. (2022); Dockhorn et al. (2022); Zhang and Chen (2022); Shaul et al. (2023); Salimans and Ho (2022); Meng et al. (2023); Luo et al. (2023); Song et al. (2023); Yin et al. (2024); Xu et al. (2024); Sauer et al. (2023); Lin et al. (2024); Schmidhuber (2020). The current SOTA method for single-step and few-step inference is Adversarial Diffusion Distillation Sauer et al. (2023), which leverages adversarial training to push the image generation to sensible images with fewer denoising steps.
@@
Fig. 9: The illustration of one time step in the learned Markov Chain in the reverse process. A deep neural network (e.g., U-Net Ronneberger et al. (2015)) learns to transform a noisy input into a less noisy image with the help of a text prompt that describes the content of the image.
@@
The diffusion model may bring to mind VAEs which encode the input data as a probability distribution and then sample from the learned latent space (Section 2.2). However, the forward process makes the diffusion model different from the VAEs as training is not required in this fixed Markov chain. More detailed information about diffusion models and its applications can be found in Yang et al. (2023).
@@ 3.1 Opening Gambit
DeepDream. Once CNNs achieved impressive results in image processing, researchers started developing visualization techniques to better understand how these neural networks see the world and perform classification. Examining each layer of a trained neural network led to the development of DeepDream Mordvintsev et al. (2015b, a), which produces surprisingly artistic images.
DeepDream generates images based on the representations learned by the neural network. It takes an input image, runs a trained CNN in reverse, and tries to maximize the activation of entire layers by applying gradient ascent Chollet (2018). DeepDream can be applied to any layer of a trained CNN. However, applying it to high-level layers is usually preferred because it provides visual patterns such as shapes or objects that are easier to recognize.
Fig. 10: (Left) A DeepDream example. Deepdream receives an input image and outputs a dreamy version in which faces and eyes emerge due to the maximization of the final layer’s activations. (Right) An illustration of neural style transfer. The content target image and style image are provided as input to the model. As output, it retains the content of the target image and applies the style of the other image.
An illustration of an original input image and its DeepDream output is shown in Fig. 10. Strikingly, the output image contains many animal faces and eyes. This is due to the fact that the original DeepDream is based on Inception Szegedy et al. (2015), which was trained using the ImageNet database Deng et al. (2009). Since there are so many examples of different dog breeds and bird species in the ImageNet database, DeepDream is biased towards those. For some people, DeepDream images resemble dream-like psychedelic experiences. In any case, although it was not its initial purpose, DeepDream inspired people to employ AI as a tool for artistic image creation.
Neural Style Transfer. A deep learning-based technique to combine the content of an image with the style of another image is called neural style transfer Gatys et al. (2015). This technique uses a pretrained CNN to transfer the style of one image to another image. A typical example is shown in Fig. 10 where the style of one image (e.g., Starry Night by Vincent Van Gogh) is applied to a content target image (e.g., a nice beach photo). Neural style transfer can be implemented by redefining the loss function in a CNN. This loss function is altered to preserve the content of the target image through the high-level activations of the CNN. At the same time, the loss function should capture the style of the other image through the activations in multiple layers. To this end, similar correlations within activations for low-level and high-level layers contribute to the loss function Chollet (2018). The result is an image that combines the content of the first input image with the style of the second input image.
@@ 3.2 Generative Adversarial Networks
ArtGAN. Tan et al. (2017) presented their model called ArtGAN, in which they show the result of a GAN trained on paintings. Although their output images looked nothing like an artwork of one of the great masters, the images seemed to capture the low-level features of artworks. This work sparked interest in the use of GANs to generate artistic images. Additionally, it challenged people to think of creativity as a concept that could be replicated by artificial intelligence.
@@
CAN. Shortly afterwards, Elgammal et al. (2017) further pushed this idea in their paper on creative adversarial neural networks (CAN). Their goal was to train a GAN to generate images that would be considered as art by the discriminator but did not fit any existing art styles. The resulting images looked mostly like abstract paintings that had a unique feel to them. Elgammal et al. (2017) then validated their work in a perceptual study where they asked human participants to rate the images on factors such as liking, novelty, ambiguity, surprisingness and complexity. In addition, they asked the participants whether they thought the image originated from a human-made painting or a computer-generated one. There were no differences in the scores on the above mentioned factors between CAN art and various abstract artworks or more traditional GAN art. However, participants more often thought that their CAN images were made by a human artist as opposed to GAN generated art.
@@
pix2pix. In 2017, Isola et al. (2017) had the innovative idea to create a conditional GAN Mirza and Osindero (2014) that receives an image as input and generates a transformed version of that image. This was achieved by training the GAN on corresponding image pairs. For example, say you have a dataset of RGB images. You can create a BW version of all these images to create image pairs (one is RGB, one is BW). What is not as trivial, however, is turning BW images into colored ones. One could manually color the images, but this is time consuming. pix2pix allows you to automate this process. The generator receives the BW version of the image pair and generates an RGB version. Next, the discriminator receives both the transformed image and the original RGB image, and has to determine which one is real and which one is fake. After the training is completed, the pix2pix GAN can transform any BW image into a colored version. The major advantage of pix2pix is that it can be applied to any dataset of image pairs without the need to adjust the training process or loss function. The same model can be used to transform sketches into paintings or BW images into colored images, simply by changing the training dataset. Many artists, as well as AI enthusiasts, have been inspired by pix2pix to create artistic images using this model (Fig. 11).
CycleGAN. Although pix2pix was a major breakthrough in generative AI, one shortcoming was that it requires corresponding image pairs for its training, which simply do not exist for all applications. For example, we do not have a corresponding photograph for every painting created by Monet. Therefore, pix2pix would not be able to turn your photograph into a Monet painting. In 2017, the same lab released CycleGAN Zhu et al. (2017), another major breakthrough in generative AI, since this GAN is able to transform your photograph into a Monet painting. CycleGAN extends pix2pix by combining two conditional GANs and ‘cycling’ between them. The first GAN’s generator might receive an image of a Monet painting and is trained to transform it to a photograph. The output image is then fed to the second generator to be transformed into a Monet painting. This transformed Monet painting and the original Monet painting are then fed to the first discriminator, whereas the photograph version of the image is compared to an existing (unpaired) photograph by the second discriminator. The same process is repeated for an existing photo by turning it into a Monet painting and back to a photo. The transformed photo is then compared to the original photo whereas the transformed Monet painting is compared to an existing (unpaired) Monet painting. In the end, the model can transform images into the other modality without having seen pairs in the training set.
@@
Fig. 11: Examples of artistic applications of pix2pix. A. Artwork “Memories of Passersby I” by Mario Klingemann. This work continuously generates a male and female looking portrait by manipulating previously generated portraits using a collection of GANs (including pix2pix). B. Screenshot from video “Learning to See: Gloomy Sunday” by Memo Akten. The original video shows a side by side of the input and output of pix2pix trained to turn ordinary video (showing household items) into artistic landscapes and flower paintings.
@@
GauGAN. In 2019, Nvidia researchers released GauGAN Park et al. (2019), named after post-impressionist painter Paul Gaugain. Similar to pix2pix, GauGAN takes an image as input and produces a transformation of that image as output. Their model uses spatially adaptive denormalization, also known as SPADE, which is a channel-wise normalization method (similar to Batch Normalization) convolved with a segmentation map. As a result, the output image can be steered with a semantic segmentation map. In addition, the input image is decoded using a VAE, which learns a latent distribution capturing the style of the image. As a result, one can generate new images of which the content is controlled by a segmentation map and the style by an existing image. Fig. 12 shows some example images generated with GauGAN.
@@
Fig. 12: Example images generated with GauGAN. One can provide a segmentation map and optionally a style reference as input. GauGAN then generates a photo-realistic version of the segmentation map in the style of the reference image. When we add a palm tree in the segmentation map, GauGAN adds a palm tree to its output in two different styles.
@@
Nvidia has since released an updated version called GauGAN2 Salian (2021). As the name suggests, it is still a GAN framework. However, this updated version can additionally perform text-to-image generation, meaning it can generate an image based on a text description as input. Earlier that year, text-to-image models became extremely popular due to the release of DALL-E Ramesh et al. (2021), a Transformer-based model which will be discussed in the next section.
@@
LAFITE. Zhou et al. (2021) proposed a GAN-based framework to perform language-free text-to-image generation, meaning they train their model solely on images. However, it is still able to generate images based on text descriptions after training is completed. They use CLIP’s Radford et al. (2021) joint semantic embedding space of text and images to generate pseudo text and image pairs. CLIP is another model that is trained to link text descriptions to the correct image and vice versa. Then, they adapt StyleGAN 2 Karras et al. (2020) to a conditional version where the text embedding is concatenated with StyleSpace, the intermediate and well-disentangled feature space of StyleGAN Karras et al. (2019).

### v3

@@ 2.4 Transformers
Convolution has become the central component for image processing applications as neural networks progressed throughout time. However, the human visual system has evolved to be sparse and efficient. Therefore, we do not process our entire visual field with the same resources, unlike convolutions. Rather, our eyes perform a fixation point strategy by means of a visual attention system which plays an essential role in human cognition [44, 45]. Inspired by the attention system in human vision, computational attention mechanisms have been developed and integrated into neural networks. One of the goals is reducing the computational burden caused by the convolution operation, while still improving the performance.
@@
These attention-based neural networks have been applied to a variety of applications such as image recognition or object tracking; see [46] for a review. In particular, a novel attention-based encoder-decoder neural network architecture was presented for neural machine translation (NMT) in 2015 [47]. The idea behind this approach is illustrated in Fig. 5, which shows how the attention mechanisms in neural networks work. In this example, the encoder takes an input sentence in English, and the decoder outputs its translation in Dutch. Both encoder and decoder include recurrent neural networks (RNNs); see [48] for more information about RNNs. Basically, the encoder outputs hidden states, and the decoder takes all the hidden states as inputs. Before processing them, the decoder applies an attention mechanism that gives each hidden state a score. Then, it multiplies each hidden state by its score to which a softmax function is applied. The weighted scores are summed, and the result leads to the context vector for the decoder. By obtaining weighted hidden states which are most associated with certain words, the decoder focuses on the relevant parts of the input during decoding.
@@
Figure 5: A neural machine translation example. The model takes an English sentence as input and translates it into Dutch. The figure shows encoder hidden states, and which words the model focuses more on (indicated by the color intensity) while translating.
@@
Attention mechanisms in neural networks have progressed rapidly, especially for NMT. One is the self-attention mechanism, which is the core building block of the Transformer [49]. The original Transformer is composed of encoder-decoder stacks which are entirely based on self-attention without any convolutional or recurrent layers. There are six identical layers in each of the encoder-decoder stacks that form the Transformer. To illustrate the model, only one encoder-decoder stack is shown in Fig. 6.
@@
Figure 6: The Transformer architecture in detail [49, 46]. (Left) The Transformer with one encoder-decoder stack. (Center) Multi-head attention. (Right) Scaled dot-product attention.
@@
The encoder-decoder stacks in the Transformer are composed of fully-connected layers and multi-head attention which is a kind of self-attention mechanism applying scaled dot-product attention within itself. As seen in Fig. 6, these attention mechanisms use three vectors for each word, namely query (Q), key (K) and value (V). These vectors are computed by multiplying the input with weight matrices $W_{q}$, $W_{k}$ and $W_{v}$ which are learned during training. In general, each value is weighted by a function of the query with the corresponding key. The output is computed as a weighted sum of the values. In the scaled dot-product attention, the dot products of the query with all keys are computed. As given in Eq. 6, each result is divided by the square root of the dimension of the keys (to have more stable gradients) and passed onto the softmax function. Finally, each softmax score is multiplied with the value [49]:
@@
Multi-head attention extends this idea by applying linear activations to the inputs (keys, values and queries) h times based on different, learned linear representations (Fig. 6). Each of the projected versions of queries, keys and values are called heads, for which the scaled dot-product attention is performed in parallel. Thus, the self-attention is calculated multiple times using different sets of query, key and value vectors. This allows the model to jointly attend to information at different positions [49]. In the last step, the projections are concatenated. Additionally, the decoder applies masked multi-head attention to ensure that only previous word embeddings (tokens) are used when predicting the next word in the sentence.
There are different Transformer architectures for various tasks [50, 51]. After the Transformer yielded major progress in natural language processing (NLP), it has been adapted to image processing tasks. Image Transformer [52] is one of these adaptations in which the Transformer is applied to image processing. Image Transformer applies self-attention in local neighborhoods for each query pixel, and performs well on image generation and image super-resolution. However, the current state-of-the-art model is Vision Transformer (ViT) [53], which splits an input image into patches before the Transformer takes the linear embeddings of these patches in sequence as input (Fig. 7). Vision Transformer performs well on image classification tasks whilst using fewer computational resources. Recent models, such as Swin transformer [54], DeiT [55], BEiT [56], and DINOv2 [57], have built on ViT and have been successful in image processing tasks.
Figure 7: The Vision Transformer [53]. In order to classify an image, it takes the input as patches, projects linearly, adds position embeddings, and uses a Transformer encoder.
@@ 2.5 Diffusion models
Today, text-to-image models such as DALL-E 3 [8] or Midjourney have turned AI into a popular tool to produce mesmerizing images. These are diffusion models which have shown great success in generating high-quality images. They have shown to outperform GANs at image synthesis [58].
@@
Unlike GANs, the training of diffusion models does not require an adversarial setting. The original Denoising Diffusion method [59] was inspired by non-equilibrium thermodynamics that systematically destroys structure in a data distribution, then restores the data. Based on this method, denoising diffusion probabilistic models or diffusion models in short, have been applied to image generation [60].
@@
Diffusion models require two main steps in the training phase (Fig. 8). At first, during the forward (diffusion) process, random noise is gradually added to the input image until the original input becomes all noise. This is performed by a fixed Markov chain which adds Gaussian noise for T successive steps. Secondly, during the reconstruction or reverse process, the model reconstructs the original data from the noise obtained in the forward process. The reverse process is defined as a Markov Chain with learned Gaussian transitions. Accordingly, the prediction of probability density at time t depends only on the state attained at time (t-1). Here, $x_{1},...,x_{T}$ are latents of the same dimensionality as the data which make the diffusion models latent variable models [60].
@@
Figure 8: The working principle of diffusion model in general [60]. Starting from an image, the forward process involves gradually adding noise (following a fixed Markov chain) until the image is all noise. In the reverse process, the original image is reconstructed step by step through a learned Markov Chain.
@@
The general structure of a diffusion model is given in Fig. 8. Estimating the probability density at an earlier time step, given the current state of the system, is non trivial, so the reverse process requires training a neural network. To this end, all previous gradients are necessary for obtaining the required estimation. For each step in the Markov Chain, a neural network learns to denoise the image. Optionally, the denoising process can be guided by text (see Section 3.3.2). In this case, a Transformer encoder maps a text prompt to tokens which are then fed to the neural network (Fig. 9). Once trained, a diffusion model can be used to generate data by simply passing random noise (and optionally a text prompt) through the learned denoising process. It should be noted that this process requires multiple denoising steps at inference time, making diffusion models slow in generating images when prompted. As a result, several methods have been proposed to speed up inference from multi-step to few-step and even single-step [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. The current SOTA method for single-step and few-step inference is Adversarial Diffusion Distillation [71], which leverages adversarial training to push the image generation to sensible images with fewer denoising steps.
@@
Figure 9: The illustration of one time step in the learned Markov Chain in the reverse process. A deep neural network (e.g., U-Net [74]) learns to transform a noisy input into a less noisy image with the help of a text prompt that describes the content of the image.
@@
The diffusion model may bring to mind VAEs which encode the input data as a probability distribution and then sample from the learned latent space (Section 2.2). However, the forward process makes the diffusion model different from the VAEs as training is not required in this fixed Markov chain. More detailed information about diffusion models and its applications can be found in [75].
@@ 3.1 Opening Gambit
DeepDream. Once CNNs achieved impressive results in image processing, researchers started developing visualization techniques to better understand how these neural networks see the world and perform classification. Examining each layer of a trained neural network led to the development of DeepDream [76, 77], which produces surprisingly artistic images.
DeepDream generates images based on the representations learned by the neural network. It takes an input image, runs a trained CNN in reverse, and tries to maximize the activation of entire layers by applying gradient ascent [26]. DeepDream can be applied to any layer of a trained CNN. However, applying it to high-level layers is usually preferred because it provides visual patterns such as shapes or objects that are easier to recognize.
An illustration of an original input image and its DeepDream output is shown in Fig. 10. Strikingly, the output image contains many animal faces and eyes. This is due to the fact that the original DeepDream is based on Inception [19], which was trained using the ImageNet database [78]. Since there are so many examples of different dog breeds and bird species in the ImageNet database, DeepDream is biased towards those. For some people, DeepDream images resemble dream-like psychedelic experiences. In any case, although it was not its initial purpose, DeepDream inspired people to employ AI as a tool for artistic image creation.
Figure 10: (Left) A DeepDream example. Deepdream receives an input image and outputs a dreamy version in which faces and eyes emerge due to the maximization of the final layer’s activations. (Right) An illustration of neural style transfer. The content target image and style image are provided as input to the model. As output, it retains the content of the target image and applies the style of the other image.
Neural Style Transfer. A deep learning-based technique to combine the content of an image with the style of another image is called neural style transfer [79]. This technique uses a pretrained CNN to transfer the style of one image to another image. A typical example is shown in Fig. 10 where the style of one image (e.g., Starry Night by Vincent Van Gogh) is applied to a content target image (e.g., a nice beach photo). Neural style transfer can be implemented by redefining the loss function in a CNN. This loss function is altered to preserve the content of the target image through the high-level activations of the CNN. At the same time, the loss function should capture the style of the other image through the activations in multiple layers. To this end, similar correlations within activations for low-level and high-level layers contribute to the loss function [26]. The result is an image that combines the content of the first input image with the style of the second input image.
@@ 3.2 Generative Adversarial Networks
ArtGAN. [80] presented their model called ArtGAN in 2017, in which they show the result of a GAN trained on paintings. Although their output images looked nothing like an artwork of one of the great masters, the images seemed to capture the low-level features of artworks. This work sparked interest in the use of GANs to generate artistic images. Additionally, it challenged people to think of creativity as a concept that could be replicated by artificial intelligence.
@@
CAN. Shortly afterwards, Elgammal et al. (2017) [81] further pushed this idea in their paper on creative adversarial neural networks (CAN). Their goal was to train a GAN to generate images that would be considered as art by the discriminator but did not fit any existing art styles. The resulting images looked mostly like abstract paintings that had a unique feel to them. They then validated their work in a perceptual study where they asked human participants to rate the images on factors such as liking, novelty, ambiguity, surprisingness and complexity. In addition, they asked the participants whether they thought the image originated from a human-made painting or a computer-generated one. There were no differences in the scores on the above mentioned factors between CAN art and various abstract artworks or more traditional GAN art. However, participants more often thought that their CAN images were made by a human artist as opposed to GAN generated art.
@@
pix2pix. In 2017, [40] had the innovative idea to create a conditional GAN [82] that receives an image as input and generates a transformed version of that image. This was achieved by training the GAN on corresponding image pairs. For example, say you have a dataset of RGB images. You can create a BW version of all these images to create image pairs (one is RGB, one is BW). What is not as trivial, however, is turning BW images into colored ones. One could manually color the images, but this is time consuming. pix2pix allows you to automate this process. The generator receives the BW version of the image pair and generates an RGB version. Next, the discriminator receives both the transformed image and the original RGB image, and has to determine which one is real and which one is fake. After the training is completed, the pix2pix GAN can transform any BW image into a colored version. The major advantage of pix2pix is that it can be applied to any dataset of image pairs without the need to adjust the training process or loss function. The same model can be used to transform sketches into paintings or BW images into colored images, simply by changing the training dataset. Many artists, as well as AI enthusiasts, have been inspired by pix2pix to create artistic images using this model (Fig. 11).
@@
Figure 11: Examples of artistic applications of pix2pix. A. Artwork “Memories of Passersby I” by Mario Klingemann. This work continuously generates a male and female looking portrait by manipulating previously generated portraits using a collection of GANs (including pix2pix). B. Screenshot from video “Learning to See: Gloomy Sunday” by Memo Akten. The original video shows a side by side of the input and output of pix2pix trained to turn ordinary video (showing household items) into artistic landscapes and flower paintings.
CycleGAN. Although pix2pix was a major breakthrough in generative AI, one shortcoming was that it requires corresponding image pairs for its training, which simply do not exist for all applications. For example, we do not have a corresponding photograph for every painting created by Monet. Therefore, pix2pix would not be able to turn your photograph into a Monet painting. In 2017, the same lab released CycleGAN [41], another major breakthrough in generative AI, since this GAN is able to transform your photograph into a Monet painting. CycleGAN extends pix2pix by combining two conditional GANs and ‘cycling’ between them. The first GAN’s generator might receive an image of a Monet painting and is trained to transform it to a photograph. The output image is then fed to the second generator to be transformed into a Monet painting. This transformed Monet painting and the original Monet painting are then fed to the first discriminator, whereas the photograph version of the image is compared to an existing (unpaired) photograph by the second discriminator. The same process is repeated for an existing photo by turning it into a Monet painting and back to a photo. The transformed photo is then compared to the original photo whereas the transformed Monet painting is compared to an existing (unpaired) Monet painting. In the end, the model can transform images into the other modality without having seen pairs in the training set.
@@
GauGAN. In 2019, Nvidia researchers released GauGAN [83], named after post-impressionist painter Paul Gaugain. Similar to pix2pix, GauGAN takes an image as input and produces a transformation of that image as output. Their model uses spatially adaptive denormalization, also known as SPADE, which is a channel-wise normalization method (similar to Batch Normalization) convolved with a segmentation map. As a result, the output image can be steered with a semantic segmentation map. In addition, the input image is decoded using a VAE, which learns a latent distribution capturing the style of the image. As a result, one can generate new images of which the content is controlled by a segmentation map and the style by an existing image. Fig. 12 shows some example images generated with GauGAN.
@@
Figure 12: Example images generated with GauGAN. One can provide a segmentation map and optionally a style reference as input. GauGAN then generates a photo-realistic version of the segmentation map in the style of the reference image. When we add a palm tree in the segmentation map, GauGAN adds a palm tree to its output in two different styles.
@@
Nvidia has since released an updated version called GauGAN2 [84]. As the name suggests, it is still a GAN framework. However, this updated version can additionally perform text-to-image generation, meaning it can generate an image based on a text description as input. Earlier that year, text-to-image models became extremely popular due to the release of DALL-E [85], a Transformer-based model which will be discussed in the next section.
@@
LAFITE. In 2021, [86] proposed a GAN-based framework to perform language-free text-to-image generation, meaning they train their model solely on images. However, it is still able to generate images based on text descriptions after training is completed. They use CLIP’s [87] joint semantic embedding space of text and images to generate pseudo text and image pairs. CLIP is another model that is trained to link text descriptions to the correct image and vice versa. Then, they adapt StyleGAN 2 [88] to a conditional version where the text embedding is concatenated with StyleSpace, the intermediate and well-disentangled feature space of StyleGAN [39].